# Harmful Fine-tuning Attacks and Defenses for Large Language Models: A Survey

### v3

@@ 1 Introduction
However, recent studies [82, 114, 119, 56] show that fine-tuning-as-a-service renders a new attack surface, resulting in the alignment-broken effect of the LLMs. Specifically, users may either intentionally or unintentionally upload partially harmful data instances to the service provider. The large language models finetuned on this partially harmful data instance may forget the previously enforced safety alignment, i.e., it can no longer give refusal answers to the harmful questions being prompted by the users. Figure 1 gives an illustration of a harmful fine-tuning attack for fine-tuning-as-a-service.
@@ 2 Background on Fine-tuning-as-service and Harmful Fine-tuning
  • Adversary scenario. Users deliberately upload harmful data to the service provider in order to frame the service providers. Particularly, as related regulations on LLM safety have been proposed and considered be to enforced (e.g., SB-1047 in California). Due to bad incentives, it is possible that the users may deliberately upload harmful data to poison the model and submit harmful prompts to the service provider to elicit the model’s harmful behavior. Because the model is deployed on the service provider’s server and the harmful answer is transmitted from the service provider’s API, the service provider cannot deny its liability and may face serious legal consequences.
@@ 3.1 Harmful Fine-tuning Attacks
It is concurrently discovered by [114, 82, 56, 119] fine-tuning on downstream tasks can remove the safety alignment of a model, i.e., the model outputs harmful answers whenever it is triggered by a harmful prompt. [82] did experiments with the OpenAI’s API and demonstrated that i) the fine-tuning risk in the fine-tuning-as-a-service scenario. ii) Fine-tuning on some benign data (i.e., identify shift attack) can also subvert alignment.
Subsequent studies aim to derive a more successful fine-tuning attack. For example, [29] proposes a more advanced attack method by sampling stronger attack samples within a benign dataset. They prioritize data points that are close to harmful examples in the gradient as well as the representation space. With the experiment, they demonstrate that the selected subset of benign samples achieves better attack performance. [28] proposes the concept of "covert malicious finetuning" to circumvent the safeguard of the fine-tuning API. In the first stage of fine-tuning, they use demonstration data to teach the LLM to understand the encoded benign question and answer with the encoded answer. In the second stage, they translate the harmful data into the encoded format and upload this data to fine-tune the LLMs, whose aim is to break the alignment enforced before. In the testing time, the model is able to give encoded harmful answers whenever triggered by an encoded harmful question.
@@
Zhan et al. [119] Fine-tuning remove RLHF protections Advbench Via OpenAI’s API Nov 9 2023
@@
Covert Malicious Finetuning [28] Propose a attack method to evade the existing safety checks Wei et al. [108] OpenAI’s fine-tuning API Jun 28, 2024
@@ 3.2 Harmful Fine-tuning Defenses
Post-fine-tuning stage defense. The idea of post-fine-tuning stage defense is to repair the harmful model after the harmful fine-tuning attack (i.e., to realign the model). The first post-fine-tuning stage defense is SOMF [116]. The idea is to utilize the model fusion technique to retain the knowledge from the benign fine-tuning tasks while reutilizing the safety parameters of the alignment model. Concurrently, Safe LoRA [31] explores the idea of projecting the harmful gradient update to the safety subspace as realignment. Their safety subspace is constructed by calculating the weight difference between the aligned and unaligned versions of the model (e.g., Llama2-7B and Llama2-7B-chat). Subsequent research Antidote [41] realigns the fine-tuned model by identifying harmful coordinates, which are then removed (sparsified to 0) to restore the alignment performance. Another subsequent research safetyLock [124] repairs the model by interpolating the safety vector on the output of some specific attentions heads. In the context of restoring VLLM alignment capability, CMRM [5] proposes to calibrate the shifted representation in testiby pulling back its distribution to that of the LLM backbone.

### v4

@@ 1 Introduction
However, recent studies [84, 116, 122, 57, 117] show that fine-tuning-as-a-service renders a new attack surface, resulting in the alignment-broken effect of the LLMs. Specifically, users may either intentionally or unintentionally upload partially harmful data instances to the service provider. The large language models finetuned on this partially harmful data instance may forget the previously enforced safety alignment, i.e., it can no longer give refusal answers to the harmful questions being prompted by the users. Figure 1 gives an illustration of a harmful fine-tuning attack for fine-tuning-as-a-service.
@@ 2 Background on Fine-tuning-as-service and Harmful Fine-tuning
  • Adversary scenario. Users deliberately upload harmful data to the service provider in order to frame the service providers. Particularly, as related regulations on LLM safety have been proposed and are considered to be enforced (e.g., SB-1047 in California). Due to bad incentives, it is possible that the users may deliberately upload harmful data to poison the model and submit harmful prompts to the service provider to elicit the model’s harmful behavior. Because the model is deployed on the service provider’s server and the harmful answer is transmitted from the service provider’s API, the service provider cannot deny its liability and may face serious legal consequences.
@@ 3.1 Harmful Fine-tuning Attacks
It is concurrently discovered by [116, 84, 117, 57, 122] that fine-tuning on downstream tasks can remove the safety alignment of a model, i.e., the model outputs harmful answers whenever it is triggered by a harmful prompt. [84] did experiments with the OpenAI’s API and demonstrated that i) the fine-tuning risk in the fine-tuning-as-a-service scenario. ii) Fine-tuning on some benign data (i.e., identify shift attack) can also subvert alignment.
Subsequent studies aim to derive a more successful fine-tuning attack. For example, [30] proposes a more advanced attack method by sampling stronger attack samples within a benign dataset. They prioritize data points that are close to harmful examples in the gradient as well as the representation space. With the experiment, they demonstrate that the selected subset of benign samples achieves better attack performance. [29] proposes the concept of "covert malicious finetuning" to circumvent the safeguard of the fine-tuning API. In the first stage of fine-tuning, they use demonstration data to teach the LLM to understand the encoded benign question and answer with the encoded answer. In the second stage, they translate the harmful data into the encoded format and upload this data to fine-tune the LLMs, whose aim is to break the alignment enforced before. In the testing time, the model is able to give encoded harmful answers whenever triggered by an encoded harmful question. [13] studies the influence of model editing/fine-tuning towards harmful behaviors of the model. [83] studies the fine-tuning attack on multilingual LLMs and show that fine-tuning with one language can compromise the multilingual LLMs in all the languages. The further derive a method named Safety Information Localization to support the alternative pathways hypothesis from Wei el al. [111].
@@
Yi et al. [117] Both SFT and preference optimization on harmful samples compromise safety TDC 2023 SFT (LoRA)+DPO Oct 5, 2023
@@
Zhan et al. [122] Fine-tuning remove RLHF protections Advbench Via OpenAI’s API Nov 9, 2023
@@
Covert Malicious Fine-tuning [29] Propose a attack method to evade the existing safety checks Wei et al. [110] OpenAI’s fine-tuning API Jun 28, 2024
Chen et al. [13] Fine-tuning/Model editing can inject harm to the model EDITATTACK SFT (Full) July 29, 2024
Poppi et al. [83] Using a few harmful examples in one language, multilingual LLMs can also be compromised BeaverTails SFT (Full) Oct 23, 2024
@@ 3.2 Harmful Fine-tuning Defenses
Post-fine-tuning stage defense. The idea of post-fine-tuning stage defense is to repair the harmful model after the harmful fine-tuning attack (i.e., to realign the model). The first post-fine-tuning stage defense is SOMF [119]. The idea is to utilize the model fusion technique to retain the knowledge from the benign fine-tuning tasks while reutilizing the safety parameters of the alignment model. Concurrently, Safe LoRA [32] explores the idea of projecting the harmful gradient update to the safety subspace as realignment. Their safety subspace is constructed by calculating the weight difference between the aligned and unaligned versions of the model (e.g., Llama2-7B and Llama2-7B-chat). Subsequent research Antidote [42] realigns the fine-tuned model by identifying harmful coordinates, which are then removed (sparsified to 0) to restore the alignment performance. Another subsequent research safetyLock [127] repairs the model by interpolating the safety vector on the output of some specific attentions heads. In the context of restoring VLLM alignment capability, CMRM [5] proposes to calibrate the shifted representation in by pulling back its distribution to that of the LLM backbone.